The sum of a number and 49 times its reciprocal is -14. Find the number.
Answer by chibisan(131)

let the number be x
its reciprocal would be 1/x

x+[49(1/x)]= -14
x+49/x = -14
x^2+49/x = -14
x^2+49 = -14x
x^2+14x+49=0
(x+7)(x+7)=0
x=-7

so the number is -7.